# SupplementaryMat

### v1

@@ Abstract
Instability Growth Timescale as an Alternative Control Parameter: Having explored the role of the activity parameter $\alpha$, which sets a characteristic timescale $\tau_{\alpha}=1/|\alpha|$ for the growth of polar order, we now turn to a complementary perspective: fixing $\alpha$ and varying the timescale associated with the growth of instabilities, defined as $\tau_{\Gamma}=\Gamma_{2}/|\Gamma_{0}|^{2}$. In previous sections, $\tau_{\Gamma}$ was kept fixed while $\alpha$ was varied to investigate its impact on the emergence of large-scale structures. However, since both parameters control the growth rate of perturbations wensink2012meso , we now ask whether varying $\tau_{\Gamma}$ at fixed $\alpha$ can produce effects analogous to varying $\alpha$ at fixed $\tau_{\Gamma}$.
@@
We fix $\alpha=-4$, below the critical threshold $\alpha_{c}$, and varied the instability growth timescale $\tau_{\Gamma}=\Gamma_{2}/|\Gamma_{0}|^{2}$ by adjusting $|\Gamma_{0}|$. For increased $\tau_{\Gamma}$ (i.e., smaller $|\Gamma_{0}|$), the energy spectra shown in Fig. 5 exhibit a power-law scaling of the form $\rm E(k)\sim k^{-3/2}$, qualitatively similar to that observed for large $|\alpha|$ ($\alpha<\alpha_{c}$ ). Conversely, reducing $\tau_{\Gamma}$ by increasing $|\Gamma_{0}|$ produces spectra resembling those at smaller $|\alpha|$,(increased $\tau_{\alpha}$). Notably, the order parameter $\phi$ sharply captures this transition, changing sign in correspondence with the shift in flow structure (Fig.5, inset).
Furthermore, to validate the generality of the order parameter, we perform two complementary simulations. First, we take the locally ordered state $\alpha=-8$ and decrease $\Gamma_{2}$ to $3.6\times 10^{-5}$, hence decreasing $\tau_{\Gamma}$ and the system goes into a disordered state with $\phi<0$. In the second simulation, we choose a disordered state $\alpha=-3$ and increased $\tau_{\Gamma}$ by increasing $\Gamma_{2}$ to $1.8\times 10^{-4}$ and the system goes into a locally ordered state with $\phi>0$ (Fig. S5). These results suggest that $\alpha$ and $\tau_{\Gamma}$ both influence the system’s instability dynamics and spectral properties, though the precise relationship between them remains to be clarified.
@@
Figure 5: The energy spectra by varying the instability growth timescale ($\tau_{\Gamma}=\frac{\Gamma_{2}}{\Gamma_{0}^{2}}$) by varying $\Gamma_{0}$ for fixed $\alpha=-4$, $\Gamma_{2}=9\times 10^{-5}$, $\beta=0.5$ and $\lambda_{0}=3.5$. Inset: The corresponding energy-based order parameter $\phi$ is plotted as a function of $\tau_{\Gamma}$, showing a sign reversal that is consistent with a transition.
Discussion: Our study uncovers a rich interplay between microscopic activity and emergent large-scale structures in the system’s flow dynamics. The results from our various analyses collectively point to a transition from disorder to a locally ordered state, as the magnitude of activity increases, through competing physical processes. Crucially, this structural transition is accompanied by the emergence of giant number fluctuations in the distribution of vortex centers, pointing towards the inhomogeneous distribution of these vortex centers in space. This structural inhomogeneity emerges from the interplay between the ordering dynamics and the inherent tendency for pattern formation. The Toner-Tu terms promote local ordering due to long-range polar alignment toner2005hydrodynamics . In contrast, the Swift-Hohenberg term introduces a counteracting influence, driving the system toward spatially chaotic structures swift1977hydrodynamic . With the increase in activity, the local polar ordering emerges and strengthens, manifesting as a marked increase in velocity correlation length. A key signature of this transition is the evolution of the velocity distribution: from unimodal Gaussian at low activity to bimodal at higher activity. Interestingly, the onset of bimodality in the velocity distribution coincides with the emergence of anomalous diffusion, as observed experimentally in bacterial turbulence by gautam2024harnessing ; ariel2015swarming and reported in simulations by mukherjee2021anomalous , suggesting a potential connection between transport behavior and changes in velocity statistics.
Notably, we find that this transition is sensitive to the strength of the higher-order derivative terms $\Gamma_{0}$ and $\Gamma_{2}$, hence $E_{\Lambda}$. This transition is robustly captured by the order parameter $\phi$, which switches sign in accordance with the onset of structural inhomogeneity. Together, these findings establish a physically grounded picture of how active stresses, alignment dynamics, and higher-order instabilities conspire to shape the multiscale organization of flows in dry active systems. The system’s transition from disorder to local order is not governed by a single control parameter but rather emerges from the nonlinear coupling of energetically distinct processes and their sensitivity to model parameters. Future investigations could extend this framework to three-dimensional active flows, where the interplay between alignment and higher-order instabilities may lead to qualitatively new regimes of spatiotemporal organization. Incorporating density fluctuations and nematic order could further elucidate the coupling between flow structure and internal degrees of freedom thampi2016active ; thampi2013velocity . On the experimental side, dense suspensions of fast-swimming bacteria or active colloids with tunable alignment and instability parameters offer promising platforms to probe the predicted transition.
Acknowledgements.
@@
We thank S. Mukherjee, S. Ray, and P. Perlekar for helpful discussions. These simulations were performed on the IITH NVIDIA V100 GPU cluster, Param Seva supercomputers through the National Supercomputing Mission, and the IITH Kanad clusters. KK acknowledges the TCS Foundation for financial support through the Research fellowship program (TCS/17/23-24/P49) KVK is grateful for funding from the Simons’ Collaboration on Wave Turbulence (award No. 651459). A. G. acknowledges SERB-DST (India) Projects MTR/2022/000232, CRG/2023/007056-G, DST (India) grant no. DST/NSM/R&D_HPC_Applications/2021/05 and grant no. SR/FST/PSI-215/2016, and IITH for Seed Grant No. IITH/2020/09 for financial support.
Supplementary Material
@@
To numerically solve Eq: 1 in main text, we have used vorticity-stream function formalism [39]. Taking curl of Eq: 1 and using from incompressibilty condition in 2d we get,
@@
where $\omega$ is ×$\nabla\times\textbf{u}$.
The simulations were performed on a square domain of size $160\times 160$ with a uniform grid resolution of $4096\times 4096$. This corresponds to a grid spacing of dx = 0.04, which sets the smallest resolvable length scale in the simulation. This grid spacing is significantly smaller than the characteristic length scale, $\Lambda=2\pi\sqrt{\frac{2\Gamma_{2}}{\Gamma_{0}}}=0.4$. It is also smaller than the typical dissipation length scale, ensuring that small-scale structures and the fine details of the flow are well captured. The choice of domain size was made to be much larger than the maximum velocity correlation length scale for the highest activity parameter $(\alpha=-9)$ used in our simulations. This ensures that the system can develop a broad range of scales without being artificially constrained by the domain boundaries and forming condensates. This setup allows the dynamics to evolve naturally, capturing both the large-scale collective motion and the finer-scale turbulent structures present in active systems. We have used the FFTW module for Fourier transforms for the pseudo-spectral method. The time marching is done using the second-order integrating factor method, where the stiff linear part is handled exactly through the integrating factor, improving stability. In the time step, $dt=2\times 10^{-4}$ for up to $2\times 10^{5}$ iterations, all the statistical averages are taken over 50 snapshots after the system reaches a steady state.
Supplementary Figure S1: (a) $V_{rms}$ increases with activity, (increasing $-\alpha$). (b)Energy spectra for different $\alpha$, showing power scaling for $\alpha<-\alpha_{c}$.
@@
The identified centroids are then used in further statistical analyses, such as computing number fluctuations across subdomains of varying size(discussed in the main text, see Fig.S3), Voronoi analysis(discussed in the next section). To analyze the distribution of intense vortices for each time step, the system is divided into subregions of area $\ell^{2}$, and estimate the number of vortex centers $n$ in each subregion, the average number of vortex centers $\bar{n}$ and the standard deviation $\Delta n$, all normalized by the total number of vortex centers. The subregion size is varied over a range with $\ell\in[5\Lambda,20\Lambda]$.
[…]
S3 Voronoi Analysis of Vortex Center Distribution
To further quantify the spatial inhomogeneity in the distribution of vortex centers, we performed a Voronoi tessellation analysis. A Voronoi tessellation divides the domain into polygonal cells such that each cell contains exactly one vortex center, and every point in a given cell is closer to its associated center than to any other. This method provides a geometric way to measure the local spacing between vortex centers and is widely used to detect clustering or depletion in point patterns[42]. By calculating the area of each Voronoi cell, we obtained a distribution of cell areas that serves as a proxy for the local density of vortex centers. In a spatially homogeneous system, Fig.S4(a), the distribution of Voronoi cell areas is expected to be narrow and peaked. However, for increasing activity, we observed a broadening of this distribution, S4(b), indicating the emergence of clustering and void-like regions, Fig.S4(c). In particular, our results show that for activity levels beyond $(\alpha_{c}=-5)$, the Voronoi area distribution departs significantly from that Gaussian distribution. This transition reflects the development of spatial inhomogeneity in the vortex center positions, consistent with our findings from the number fluctuation analysis. The inset displays the interquartile range (IQR), which serves as a measure of the spread in the area distribution. The IQR is defined as the difference between the 75th percentile $(Q_{3})$ and the 25th percentile $(Q_{1})$ of the distribution:
@@
The IQR effectively captures the range within which the central $50\%$ of the data fall. The IQR of area distribution increases with activity, indicating that the distribution of Voronoi areas becomes more spread out at higher activity levels.
@@
Supplementary Figure S4: (a) and (b) represent the area constructed by Voronoi tessellation based on the filtered vortex index for $\alpha=-1$ and $\alpha=-8$, respectively. (e) The probability distribution of the areas of these Voronoi polygons. The inset shows the interquartile range (IQR) of the distribution
S4 Order Parameter
@@
This competition governs the emergence and breakdown of mesoscale structures. For instance, at a fixed activity level $\alpha=-3$, increasing $\Gamma_{2}$ leads to a transition from a disordered to an ordered state, as reflected by a change in the sign of the order parameter $\phi$ and a corresponding shift in the velocity distribution from unimodal to bimodal, Fig.S5(b). Similarly, at $\alpha=-8$, decreasing $\Gamma_{2}$ results in a breakdown of local order, with $\phi$ switching from positive to negative and the velocity statistics reverting to a unimodal form, Fig.S5(c).
@@
Supplementary Figure S5: (a) Order parameter $\phi$ as a function of activity $\alpha$, showing the emergence of new data points obtained by varying the coefficient $\Gamma_{2}$. The additional points reveal a shift in the critical activity where the transition in flow structure occurs, compared to the fixed-parameter case. (b) Velocity magnitude distribution for $\alpha=-3$, illustrating a transition from an unimodal Gaussian profile to a bimodal distribution as $\phi$ becomes positive, indicating the onset of local ordering. (c) Corresponding distribution for $\alpha=-8$, where the sign reversal in $\phi$ leads to a collapse of bimodality and a return to Gaussian statistics, signaling the loss of extended flow coherence.

### v2

@@ Abstract
Instability growth timescale as an alternative control parameter— Having explored the role of the activity parameter $\alpha$, which sets a characteristic timescale $\tau_{\alpha}=1/|\alpha|$ for the growth of polar order, we now turn to a complementary perspective: fixing $\alpha$ and varying the timescale associated with the growth of instabilities, $\tau_{\Gamma}$. In previous sections, $\tau_{\Gamma}$ was kept fixed while $\alpha$ was varied to investigate its impact on the emergence of large-scale structures. However, since both parameters control the growth rate of perturbations [48], we now ask whether varying $\tau_{\Gamma}$ at fixed $\alpha$ can produce effects analogous to varying $\alpha$ at fixed $\tau_{\Gamma}$.
@@
We fix $\alpha=-4$, below the critical threshold $\alpha_{c}$, and vary the instability growth timescale, $\tau_{\Gamma}$ by adjusting $|\Gamma_{0}|$. For increased $\tau_{\Gamma}$ (i.e., smaller $|\Gamma_{0}|$), the energy spectra shown in Fig. 3 exhibit a power-law scaling of the form $\rm E(k)\sim k^{-3/2}$, qualitatively similar to that observed for large $|\alpha|$  [27] (Fig S1(b)) . Conversely, reducing $\tau_{\Gamma}$ by increasing $|\Gamma_{0}|$ produces spectra resembling those at smaller $|\alpha|$, (increased $\tau_{\alpha}$). Notably, the order parameter $\phi$ sharply captures this transition, changing sign in correspondence with the shift in flow structure (Fig.3, inset).
To further assess the generality of the order parameter, we extended the analysis to two additional representative values of $\tau_{\Gamma}$ ($\Gamma_{2}$), as shown in Fig. 2(c). For decreased $\tau_{\Gamma}$ (corresponding to $\Gamma_{2}=3.6\times 10^{-5}$), the system remains disordered with $\phi<0$ for all values of $\alpha$ considered, and no local ordering emerges. In contrast, for an increased $\tau_{\Gamma}$ (larger $\Gamma_{2}=1.8\times 10^{-4}$), the system undergoes a transition to a locally ordered state with $\phi>0$ already at much lower activity levels ($\alpha=-2$). These results demonstrate that both $\alpha$ and $\tau_{\Gamma}$ act as key control parameters for the system’s dynamics, jointly determining the onset of local ordering. Correspondingly, the energy spectra exhibit distinct signatures, reflecting the redistribution of energy across scales as the system transitions between disordered and locally ordered states (Fig. S8(c)). In parallel, the velocity statistics undergo a qualitative change: while the distribution remains unimodal in the disordered regime, increasing $\tau_{\Gamma}$ drives the system into a locally ordered state characterized by a bimodal velocity distribution (Fig. S8(b)).
@@
Figure 3: The energy spectra by varying the instability growth timescale $\tau_{\Gamma}$ by varying $\Gamma_{0}$ for fixed $\alpha=-4$, $\Gamma_{2}=9\times 10^{-5}$, $\beta=0.5$ and $\lambda_{0}=3.5$. Inset: The corresponding energy-based order parameter $\phi$ is plotted as a function of $\tau_{\Gamma}$, showing a sign reversal that is consistent with a transition.
Discussion— We identify a previously unrecognized structural transition in active turbulence, marked by the simultaneous onset of local ordering and giant number fluctuations. The Toner-Tu terms promote local ordering due to long-range polar alignment [44, 43]. In contrast, the Swift-Hohenberg term introduces a counteracting influence, driving the system toward spatially chaotic structures [42]. Increasing activity enhances local polar order and consequently increases the velocity correlation length. In the high-activity regime, signatures of local ordering are reflected in the single-point velocity statistics, which become bimodal. This regime also coincides with the onset of anomalous diffusion in bacterial turbulence, reported in simulations [26] and observed experimentally [2, 10]. This suggests a potential connection between transport behavior and the underlying velocity statistics. This transition is sensitive to the strength of the higher-order derivative terms $\Gamma_{0}$ and $\Gamma_{2}$, hence $E_{\Lambda}$ and is robustly captured by the order parameter $\phi$, which switches sign in accordance with the onset of structural inhomogeneity. Our analysis reveals that the velocity field transitions into a state in which ordered vortical domains, characterized by large correlations, interact with disordered fluctuations, leading to the observed universal behavior at high activity. Beyond serving as a quantitative marker, the energy-based order parameter encapsulates this competition among active driving, instability, and kinetic dissipation, providing a unified energetic criterion for pattern selection in active turbulence. We anticipate that this framework will help connect statistical universality to structural organization in other active and driven dissipative systems [25]. The predicted transition could be probed in dense suspensions of fast-swimming bacteria or active colloids with tunable alignment by measuring the emergence of bimodal velocity statistics and correlated vortex clustering. The recent 3D simulations of the Toner-Tu-Swift-Hohenberg model reveal stable flocking regimes at high activity that are absent in 2D [33], suggesting that the coexistence of ordered and chaotic states identified here may acquire a qualitatively new character in three-dimensional active systems. Together, these findings provide a unified framework linking structural, statistical, and energetic signatures of turbulence in active matter.
@@
Acknowledgements — We thank S. Mukherjee, S.S. Ray, P. Perlekar, S. Musacchio, and S. Islam for helpful discussions. These simulations were performed on the IITH NVIDIA V100 GPU cluster, Param Seva supercomputers through the National Supercomputing Mission, and the IITH Kanad clusters. KK acknowledges the TCS Foundation for financial support through the Research fellowship program (TCS/17/23-24/P49). KVK is grateful for funding from the Simons’ Collaboration on Wave Turbulence (award No. 651471). A. G. acknowledges SERB-DST (India) Projects MTR/2022/000232, CRG/2023/007056-G, DST (India) grant no. DST/NSM/R&D_HPC_Applications/2021/05 and grant no. SR/FST/PSI-215/2016, and IITH for Seed Grant No. IITH/2020/09 for financial support.
Supplemental Material
@@
To numerically solve Equation (1) in the main text, we have used vorticity-stream function formalism [39]. Taking curl of Equation (1) and using from incompressibilty condition in 2d we get,
@@
where $\omega$ is $\nabla\times\textbf{u}$.
The simulations were performed on a square domain of size $160\times 160$ with a uniform grid resolution of $4096\times 4096$. This corresponds to a grid spacing of dx = 0.04, which sets the smallest resolvable length scale in the simulation. This grid spacing is significantly smaller than the characteristic length scale, $\Lambda=2\pi\sqrt{\frac{2\Gamma_{2}}{|\Gamma_{0}|}}=0.4$. It is also smaller than the typical dissipation length scale, ensuring that small-scale structures and the fine details of the flow are well captured. The choice of domain size was made to be much larger than the maximum velocity correlation length scale for the highest activity parameter $(\alpha=-9)$ used in our simulations. This ensures that the system can develop a broad range of scales without being artificially constrained by the domain boundaries and forming condensates. This setup allows the dynamics to evolve naturally, capturing both the large-scale collective motion and the finer-scale turbulent structures present in active systems. We have investigated the system for $\alpha\in[-9,-1]$. The root-mean-square velocity, $u_{\mathrm{rms}}$, is shown in Fig. S1(a). It increases linearly with $\alpha$ for $\alpha>\alpha_{c}$ [4], but deviates from this linear trend for $\alpha<\alpha_{c}$, where $\alpha_{c}=-5$. The corresponding spectra are shown in Fig. S1(b), where the 3/2-3/2 slope emerges for $\alpha<\alpha_{c}$ [27]. We have used the FFTW module for Fourier transforms for the pseudo-spectral method. The time marching is done using the second-order integrating factor method, where the stiff linear part is handled exactly through the integrating factor, improving stability. The simulations use a time step of $dt=2\times 10^{-4}$ for up to $2\times 10^{5}$ iterations, corresponding to a total duration of $100\tau_{\Gamma}$, where $\tau_{\Gamma}=2\pi\sqrt{2}\frac{\Gamma_{2}}{|\Gamma_{0}|^{2}}$ is the instability growth timescale. Statistical averages are computed over the interval $50\tau_{\Gamma},100\tau_{\Gamma}]$ using 50 snapshots.
Experiments on B. subtilis, at normal oxygen concentration, report mean swimming speeds of $\simeq 25~\mu\mathrm{m}/\mathrm{s}$ with a field of view of approximately $400~\mu\mathrm{m}\times 400~\mu\mathrm{m}$  [48, 4, 16]. We map these experimental values onto the constant velocity scale $v_{c}$ and the simulation domain size $L\times L$, respectively. This mapping results in scale factors of $25/v_{c}$ for velocities and $4\times 10^{-2}/L$ for lengths when relating our parameters to experimental observations. With this correspondence, the characteristic velocities $_{c}, vv$, and $u_{\mathrm{rms}}$ in the simulations correspond to $\simeq 25~\mu\mathrm{m}/\mathrm{s}$, $\simeq(25~\mu\mathrm{m}/\mathrm{s}$ to $75~\mu\mathrm{m}/\mathrm{s})$, and $\simeq(4.5~\mu\mathrm{m}/\mathrm{s}$ to $70~\mu\mathrm{m}/\mathrm{s})$, respectively. Likewise, the parameter $\Lambda$, which sets the characteristic linear size of vortical structures, corresponds to $\simeq 25~\mu\mathrm{m}$.
Supplementary Figure S1: (a) $u_{rms}$ increases with activity, (increasing $\alpha$). (b)Energy spectra for different $\alpha$, showing power scaling for $\alpha<\alpha_{c}$.
@@
The identified centroids are then used in further statistical analyses, such as computing number fluctuations across subdomains of varying size(discussed in the main text, see Fig. S3), Voronoi analysis (discussed in the next section). To analyze the distribution of intense vortices for each time step, the system is divided into subregions of area $\ell^{2}$, and estimate the number of vortex centers $n$ in each subregion, the mean number of vortex centers $\bar{n}$ and the standard deviation $\Delta n$, all normalized by the total number of vortex centers. The subregion size is varied over a range with $\ell\in[5\Lambda,20\Lambda]$. In Fig. S4, we plot the probability distribution function (PDF) of $\rm n$ around for different $\alpha$, in a given subregion. We note that for decreasing $\alpha$, the $P(\frac{\rm n}{\bar{n}})$ broadens, indicating stronger fluctuations in the number of vortex centers. The inset displays the interquartile range (IQR), which serves as a measure of the spread in the number distribution. The IQR is defined as the difference between the 75th percentile $(Q_{3})$ and the 25th percentile $(Q_{1})$ of the distribution:
@@
The IQR effectively captures the range within which the central $50\%$ of the data fall.
Supplementary Figure S3: (a) and (e) show the vorticity field at a representative zoomed-in snapshot for weak ($\alpha=-1$) and strong ($\alpha=-6$) activity, respectively, revealing a transition from spatially disordered vortex structures to highly structured patterns with dense vortical regions and zero or less vorticity regions. (b) and (f) display the corresponding extracted vortex centers, obtained by identifying regions below a threshold in the Okubo–Weiss parameter and computing their intensity-weighted centroids. (c) and (g) show the probability distribution functions of the number of vortex centers $n_{\ell}$ within subdomains of varying size $\ell$, with[$\ell\in[5\Lambda,20\Lambda$], highlighting the change in spatial homogeneity. (d) and (h) present the number fluctuations $\Delta n$ as a function of the average $\bar{n}_{\ell}$, respectively, for the two activity levels. For low activity ($\alpha=-1$), $\Delta n\sim\bar{n}_{\ell}^{0.5}$, indicating homogeneous distribution, while for high activity ($\alpha=-6$), the scaling approaches $\Delta n\sim\bar{n}_{\ell}^{0.8}$, consistent with inhomogenity in vortex centers distribution.
Supplementary Figure S4: Probability distribution functions (PDF) of the normalized vortex count $(\rm n/\bar{n})$ in subdomains of fixed linear size $\ell$ for different activity values. As $|\alpha|$ increases the distributions broaden and their maxima shift from $\rm n/\bar{n}\simeq 1$ toward lower values, indicating that typical subdomains contain fewer vortex centers, while strong fluctuations become more likely. The inset shows the interquartile range (IQR) of the distribution.
S3 Static Structure factor
We compute the static structure factor of the number density of vortex centers, which characterizes equal-time spatial correlations in number density. To this end, the domain is coarse-grained into square boxes of size $l=5\Lambda$. The local density fluctuation is defined as $\delta\rho_{l}=\rho_{l}-\rho_{0}$, where $\rho_{l}=n/l^{2}$, $n$ is the number of vortex centers in a given box of size $l^{2}$ and $\rho_{0}$ is the mean density. The Fourier transform $\delta\rho_{l}(\mathbf{k})$ is then used to compute the structure factor according to the standard definition [43, 9].
$S(k)=\frac{1}{N}\left\langle|\delta\rho_{l}(\mathbf{k})|^{2}\right\rangle.$ (S4)
For large activities ($\alpha<\alpha_{c}$), the resulting structure factor $S(k)$ increases with decreasing wave number $k$ at intermediate scales, reaches a maximum at a finite $k$, and subsequently decreases as $k\to 0$. Consequently, $S(k)$ remains finite in the long-wavelength limit, in contrast to the divergence observed in MIPS [9]. This behavior indicates that giant number fluctuations are present only over short to intermediate length scales (at mesoscale) and are suppressed at larger scales as shown for $\alpha=-6$ and 8-8 in Fig S5(a). For smaller activities ($\alpha>\alpha_{c}$), where giant number fluctuations are absent, the structure factor does not have a characteristic peak and saturates to a constant value as $k\to 0$, shown for $\alpha=-1$ in Fig S5(a). Furthermore, for a given value of $\alpha$ for example ($\alpha=-6$), the characteristic peak of the structure factor (seen at $\sim 50\Lambda$) is unaffected by changes in system size $L$ (we vary system size between $100\Lambda-500\Lambda$, Fig S5(b)).
Supplementary Figure S5: (a) Static Structure factor for different activity. (b) Static Structure factor for fixed activity ($\alpha=-6$) and varying system size L.
S4 Voronoi Analysis of Vortex Center Distribution
To further quantify the spatial inhomogeneity in the distribution of vortex centers, we performed a Voronoi tessellation analysis. A Voronoi tessellation divides the domain into polygonal cells such that each cell contains exactly one vortex center, and every point in a given cell is closer to its associated center than to any other. This method provides a geometric way to measure the local spacing between vortex centers and is widely used to detect clustering or depletion in point patterns [30]. By calculating the area of each Voronoi cell, we obtained a distribution of cell areas that serves as a proxy for the local density of vortex centers. In a spatially homogeneous system, Fig. S6(a), the distribution of Voronoi cell areas is expected to be narrow and peaked. However, for increasing activity, we observed a broadening of this distribution, Fig. S6(b), indicating the emergence of clustering and void-like regions, Fig. S6(c). In particular, our results show that for activity levels below $(\alpha_{c}=-5)$, the Voronoi area distribution departs significantly from that Gaussian distribution. This transition reflects the development of spatial inhomogeneity in the vortex center positions, consistent with our findings from the number fluctuation analysis. The inset shows the IQR of area distribution increases with activity, indicating that the distribution of Voronoi areas becomes more spread out at higher activity levels.
@@
Supplementary Figure S6: (a) and (b) represent the area constructed by Voronoi tessellation based on the filtered vortex index for $\alpha=-1$ and $\alpha=-8$, respectively. (c) The probability distribution of the area normalised by $\Lambda^{2}$ of these Voronoi polygons. The inset shows the interquartile range (IQR) of the distribution.
S5 Non-dimensionalization and polar reduction for an isolated vortex
S5.1 Dimensional model
We begin with Toner–Tu–Swift–Hohenberg (TTSH) type equation for an incompressible active fluid,
$\partial_{t}\mathbf{u}+\lambda_{0}\,(\mathbf{u}\!\cdot\!\nabla)\mathbf{u}=-\nabla p\;+\;\Gamma_{0}\nabla^{2}\mathbf{u}\;-\;\Gamma_{2}\nabla^{4}\mathbf{u}\;-\;\bigl(\alpha+\beta|\mathbf{u}|^{2}\bigr)\mathbf{u}$ (S5)
S5.2 Choice of scales and non-dimensionalization
Scaling length by $\Lambda=\sqrt{\frac{\Gamma_{2}}{|\Gamma_{0}|}}$, time with $\tau_{\alpha}=\frac{1}{|\alpha|}$, velocity with $v=\sqrt{\frac{-\alpha}{\beta}}$, and pressure accordingly:
$\mathbf{x}=\Lambda\,\mathbf{x}^{\prime},\qquad t=\frac{t^{\prime}}{\alpha},\qquad\mathbf{u}=v\,\mathbf{u}^{\prime},\qquad p=\alpha\Lambda v\,p^{\prime}\qquad$
Under these, derivatives rescale as
$\nabla=\frac{1}{\Lambda}\nabla^{\prime},\qquad\partial_{t}=\alpha\,\partial_{t^{\prime}}$
Substituting into (S5) and dividing by $\alpha v$, gives
$\partial_{t^{\prime}}\mathbf{u}^{\prime}+\frac{\lambda_{0}v}{\alpha\Lambda}\,(\mathbf{u}^{\prime}\!\cdot\!\nabla^{\prime})\mathbf{u}^{\prime}=-\nabla^{\prime}p^{\prime}+\underbrace{\frac{\Gamma_{0}}{\alpha\Lambda^{2}}}_{=\,R}\nabla^{\prime 2}\mathbf{u}^{\prime}-\underbrace{\frac{\Gamma_{2}}{\alpha\Lambda^{4}}}_{=\,R}\nabla^{\prime 4}\mathbf{u}^{\prime}-\Bigl(1-|\mathbf{u}^{\prime}|^{2}\Bigr)\mathbf{u}^{\prime}.$ (S6)
where,
$R\;=\;\frac{\Gamma_{0}}{\alpha\Lambda^{2}}=\frac{\Gamma_{2}}{\alpha\Lambda^{4}}=\frac{\Gamma_{0}^{2}}{\alpha\Gamma_{2}}$
and dropping primes, the non-dimensional system is
$\partial_{t}\mathbf{u}+\frac{\lambda_{0}v}{\alpha\Lambda}\,(\mathbf{u}\!\cdot\!\nabla)\mathbf{u}=-\nabla p\;+\;R\,\bigl(\nabla^{2}-\nabla^{4}\bigr)\mathbf{u}\;-\;\bigl(1-|\mathbf{u}|^{2}\bigr)\mathbf{u}.$ (S7)
S5.3 Steady azimuthal vortex ansatz in polar coordinates
We consider a steady, purely azimuthal, axisymmetric vortex in 2D polar coordinates $(r,\theta)$:
$\partial_{t}\mathbf{u}=0,\qquad\mathbf{u}(r)=u_{r}\,\hat{\mathbf{r}}+u_{\theta}\,\hat{\bm{\theta}},\qquad u_{r}(r)=0,\quad u_{\theta}(r)=f(r).$
With $p=p(r)$ by symmetry, the $\theta$–component of the advective term and of $\nabla p$ vanish. The vector Laplacian acting on the $\theta$–component (no $\theta$-dependence) reads
$(\nabla^{2}\mathbf{u})_{\theta}\;=\;\mathcal{L}[f]:=f^{\prime\prime}+\frac{1}{r}f^{\prime}-\frac{f}{r^{2}}.$ (S8)
Applying the same (componentwise) Laplacian again gives the bilaplacian contribution
$(\nabla^{4}\mathbf{u})_{\theta}\;=\;\mathcal{L}\!\bigl[\mathcal{L}[f]\bigr]=f^{(4)}+\frac{2}{r}f^{(3)}-\frac{3}{r^{2}}f^{\prime\prime}+\frac{3}{r^{3}}f^{\prime}-\frac{3}{r^{4}}f.$ (S9)
(Equations (S8)–(S9) follow by straightforward differentiation of $\mathcal{L}$.)
S5.4 Reduced fourth-order ODE for the vortex profile
Taking the $\theta$–component of (S7) under the above ansatz yields
$R\,\bigl(\mathcal{L}^{2}f-\mathcal{L}f\bigr)\;=\;-(1-f^{2})\,f.$ (S10)
Expanding with (S8)–(S9) gives the explicit radial ODE
$R\left[f^{(4)}+\frac{2}{r}f^{(3)}-\left(1+\frac{3}{r^{2}}\right)f^{\prime\prime}+\left(\frac{3}{r^{3}}-\frac{1}{r}\right)f^{\prime}+\left(\frac{1}{r^{2}}-\frac{3}{r^{4}}\right)f\right]\;=\;-(1-f^{2})\,f.$ (S11)
S5.5 Small-$r$ asymptotics
For $r\ll\Lambda$, we assume a power-law ansatz for the azimuthal velocity,
$f(r)\sim r^{a}.$ (S12)
Substituting this form into Eq. S11 and retaining the dominant $\mathcal{O}\!\left(\tfrac{1}{r^{4}}\right)$ contributions yields an algebraic equation for $a$.
$(a-1)^{2}(a-3)(a+1)=0$ (S13)
The distinct roots are
$a=1,\qquad a=3,\qquad a=-1$ (S14)
The negative exponents is discarded since it leads to divergences in $f(r)$. The cubic branch is regular but subdominant, so the small-$r$ behaviour is dominated by the linear solution
$f(r)\sim Ar.$
Supplementary Figure S7: Radial ($u_{r}$) and azimuthal ($u_{\theta}$) velocity components near an isolated vortex point as a function of radius $r$. The profiles show that the velocity is primarily aligned in the azimuthal direction with negligible radial component. For $r>\Lambda$ (asymptotic limit), the azimuthal velocity converges to $v=\sqrt{|\alpha|/\beta}$.
S6 Order Parameter
@@
This competition governs the emergence and breakdown of mesoscale structures. For example, at a fixed activity level of $\alpha=-3$, increasing $\Gamma_{\tau}$ from 0.40 to 0.80 by increasing $\Gamma_{2}$, induces a transition from a disordered to a locally ordered state, as indicated by the change in the sign of the order parameter ($\phi<0\to\phi>0$) and a corresponding shift in the velocity distribution from unimodal to bimodal Fig. S8(b). Conversely, at $\alpha=-8$, decreasing $\Gamma_{\tau}$ to 0.20 from 0.40 by decreasing $\Gamma_{2}$, destabilizes the ordered state, leading to $\phi$ switching from positive to negative and the velocity statistics reverting to a unimodal form Fig. S8(b). Correspondingly, the energy spectra exhibit distinct changes, reflecting the redistribution of energy across scales as the system transitions between ordered and disordered regimes, Fig. S8(c).
@@
Supplementary Figure S8: ((a) Order parameter $\phi$ as a function of activity $\alpha$ for three different values of $\tau_{\Gamma}$ obtained by varying $\Gamma_{2}$. The curves highlight how the critical activity at which the transition in flow structure occurs shifts with $\tau_{\Gamma}$. (b, top) Probability distribution function of normalized velocity component, for $\alpha=-3$, demonstrating a transition from a unimodal Gaussian profile to a bimodal form as $\phi$ becomes positive, marking the onset of local ordering. The points highlighted with black borders indicate the cases for which the velocity component distributions are shown in panel (b). (b, bottom) Corresponding normalized velocity component distribution for $\alpha=-8$, illustrating the reverse transition from a bimodal to a Gaussian profile as local order breaks down. (d) Energy spectra for the two representative cases: at $\alpha=-8$, the spectrum changes from a $k^{-3/2}$ power law to no clear scaling as the system enters a disordered state; conversely, at $\alpha=-3$, a $k^{-3/2}$ scaling emerges from an initially structureless spectrum, reflecting the onset of ordering.